OpenStudy (anonymous):

Multiply and simplify the radical if possible.. Just need someone to point out the steps =)

OpenStudy (anonymous):

\[\sqrt{3x}(\sqrt{x} - 3 \sqrt{3})\]

OpenStudy (anonymous):

Well i know i need to work inside the parenthesis first..